OpenStudy (anonymous):

The HCF of 27xy, 36x, 9y

OpenStudy (anonymous):

27xy = 3*3*3*x*y right?

OpenStudy (anonymous):

That's is right. What about for the whole together

OpenStudy (anonymous):

Similarly can u express 36x in that form

OpenStudy (anonymous):

Yeah. 36 is 2*2*3*3*x

OpenStudy (anonymous):

good

OpenStudy (anonymous):

and 9y= ?

OpenStudy (anonymous):

The problem is not for the individual numbers but for them together

OpenStudy (anonymous):

9y is 3*3*y

OpenStudy (anonymous):

27xy = 3*3*3*x*y 36x=2*2*3*3*x 9y=3*3*y Now, HCF= 3*3

OpenStudy (anonymous):

Just take numbers that are common in all the above

OpenStudy (anonymous):

exressions

OpenStudy (anonymous):

Thanks . I get it now

OpenStudy (anonymous):

welcome